OpenStudy (summersnow8):

Its rate of chirping is described by the following formula: N= number of chirps per13.0 seconds T−40.0 In this expression, T is the temperature in degrees Fahrenheit. If the temperature is 42 degrees Fahrenheit, how long does it take for the cricket to chirp 6 times? 6s? 42s? 39s? 12s?

OpenStudy (whpalmer4):

Funny, I was just today reading a book which uses this very same example! N = T - 40.0 gives the number of chirps in 13.0 seconds. at T = 42 degrees F, N = 42 - 40 = 2 chirps in 13.0 seconds. So our chirping rate at T = 42.0 degrees is 2 chirps/13.0 seconds. To get the time needed for 6 chirps, we do 6 chirps * 13.0 seconds/2 chirps = 78.0/2 second = 39.0 seconds. This makes sense, as 39 seconds is 3 * 13 seconds, so we get 3*2 = 6 chirps.

OpenStudy (summersnow8):

nice! I will see if I can figure out the others :)